OpenStudy (anonymous):

What are the factors of x2 + 7x + 12? @dpasingh @Algebra2013

OpenStudy (anonymous):

\[x^2 + 7x + 12\]\[(x+3) (x+4)\]

OpenStudy (anonymous):

Thank you lots.a

OpenStudy (anonymous):

What two numbers WHEN ADDED equal 7 and WHEN MULTIPLIED equal 12? Thats the answer..

OpenStudy (anonymous):

\[ x^2 + 7x + 12= x^2 + 4x+3X + 12 =X(X+4)+3(X+4)= (X+3)(X+4)\]

OpenStudy (anonymous):

Your welcome :)

OpenStudy (anonymous):

Oh that makes more sense now!